# Xing Huang(a), Eva Kopfer(b), Panpan Ren(c) a)

### v3

@@ 2. $(2)$
(1)(i) Taking $y={\bf 0}$ in ((H)), we get
So, it follows from (2.15) and (2.2) that
for some constant $C_{0}>0$. Whence, an application of [44, Lemma 2.7], along with (2.15) and (2.23), yields that (2.13) has a unique IPM, written as $\Phi(\mu).$ Next, once $K_{2}>K_{I}$ is valid, [44, Lemma 2.8] implies that, for some constant $M>0,$ $\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$ is an invariant set of $\Phi$ so the second assertion in (1) holds true.
(ii) Itô’s formula, (2.15) as well as (2.23) imply that for $\varepsilon=\frac{K_{2}}{4\delta_{1}}$ and $\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}),$
$\text{\rm{d}}\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}$ $\leq\varepsilon\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}\left\{-K_{2}|\tilde{X}_{t}^{\mu}|^{2}+2|\tilde{X}_{t}^{\mu}||b({\bf 0},\delta_{\bf 0})|+K_{I}\mu(|\cdot|^{2})+C_{0}+2\varepsilon\delta_{1}|\tilde{X}_{t}^{\mu}|^{2}\right\}\text{\rm{d}}t$
$\leq$ $\varepsilon\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}\left\{-\frac{K_{2}}{4}|\tilde{X}_{t}^{\mu}|^{2}+\frac{4|b({\bf 0},\delta_{\bf 0})|^{2}}{K_{2}}+K_{I}M+C_{0}\right\}\text{\rm{d}}t$
$\varepsilon\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}\left\{-\frac{K_{2}}{4}|\tilde{X}_{t}^{\mu}|^{2}+\frac{4|b({\bf 0},\delta_{\bf 0})|^{2}}{K_{2}}+K_{I}M+C_{0}\right\}\leq c_{1}(\varepsilon)-c_{2}(\varepsilon)\text{\rm{e}}^{\varepsilon|\tilde{X}_{t}^{\mu}|^{2}}.$
Substituting this into (2.2) and taking $\tilde{X}_{0}^{\mu}={\bf 0}$ enables us to derive that for some constant $C_{1}>0$
This obviously implies (2.20) by a tightness argument.
(iii) By [42, Lemma 2.6], there exist constants $p>1$ depending on $\delta_{1},\delta_{2}$, $h_{1}>0$ depending on $K_{1},r_{0}$, $\delta_{1},\delta_{2}$ and $h_{2}>0$ depending on $\delta_{1},\delta_{2}$ such that the dimension free Harnack inequality holds, i.e. for any $t>0$,
$(\tilde{P}_{t}^{\mu}|f|)^{p}(x)$ $\leq(\tilde{P}_{t}^{\mu}|f|^{p})(y)$
(2.25) ×$\qquad\quad\times\exp\left\{h_{1}t+\frac{h_{2}K_{2}|x-y|^{2}}{\text{\rm{e}}^{K_{2}t}-1}\right\},\ \ \mu\in\mathscr{P}_{2}(\mathbb{R}^{d}),x,y\in\mathbb{R}^{d},f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
When $\sigma=I_{d\times d}$, the Harnack inequality above has been established in [41, p.1138]. By following the strategy adopted in [40, Corollary 1.3], it is more or less standard to derive from (2.2) that for some locally bounded function $(\infty,0)\ni t\mapsto C_{t}\in[0,\infty)$,
$\sup_{\{\Phi(\mu)\}(|f|^{p})\leq 1}(\tilde{P}_{t}^{\mu}|f|)^{2p}(x)$ $\leq C_{t}\exp\left\{\frac{h_{2}K_{2}}{\text{\rm{e}}^{K_{2}t}-1}|x|^{2}\right\},\ \ \mu\in\mathscr{P}_{2}(\mathbb{R}^{d}),x\in\mathbb{R}^{d}.$
(2.26) $\|\tilde{P}_{t_{0}}^{\mu}\|_{L^{p}(\Phi(\mu))\rightarrow L^{2p}(\Phi(\mu))}^{2p}\leq C_{t_{0}}\{\Phi(\mu)\}(\text{\rm{e}}^{\varepsilon|\cdot|^{2}})\leq C_{t_{0}}C_{1},\ \ \mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d}).$
(2) If $\sigma(x,\gamma)=\sigma(\gamma)$, by (2.4) with $s=0$, we have
@@
Since $\{\Phi(\mu):\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})\}$ is tight due to (2.20), according to Lemma 2.5(2), in order to derive the desired log-Sobolev inequality (2.21), it remains to prove that for any $R>0$ and $M>0$, there exist positive constants $C_{1}$ and $C_{2}$ depending on $R$, $M$ such that
@@
for $\rho_{\mu}:=\frac{\text{\rm{d}}\Phi(\mu)}{\text{\rm{d}}x}$. Since $b(x,\mu)$ is locally bounded in $(x,\mu)\in\mathbb{R}^{d}\times\mathscr{P}_{2}(\mathbb{R}^{d})$, for any $R>0,M>0$, there exists $\eta(R,M)>0$ such that
@@
Noting $(\sigma\sigma^{\ast})(\mu)\geq\delta_{2}I_{d\times d}$, by [5, Corollary 1.2.11], for any $R>0$ and $M>0$, there exists a constant $C_{0}$ depending on $R$, $\eta(R,M)$ and $\delta_{2}$ such that
@@
2.3 Exponential ergodicity of (2.12) in $\mathbb{W}_{2}$ and relative entropy under partially dissipative condition
@@
Assume (H) and that $\sigma(x,\gamma)=\sigma(\gamma)$. If $K_{I}$ is small enough, then (2.12) has a unique IPM $\mu_{\infty}$ which satisfies the log-Sobolev inequality
(2.29) $\mu_{\infty}(f^{2}\log f^{2})\leq C_{L}\mu_{\infty}(|\nabla f|^{2}),\ \ \mu_{\infty}(f^{2})=1,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d})$
for some constant $C_{L}>0$. Moreover, there exist constants $c>0,\lambda>0$ such that
@@
(2.32) $\mathrm{Ent}(P_{t}^{\ast}\mu_{0}|\mu_{\infty})\leq\bar{c}^{2}\text{\rm{e}}^{-2\lambda t}\mathrm{Ent}(\mu_{0}|\mu_{\infty}),\ \ t\geq 1,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{d})$
@@
(1) Recall $\Phi(\mu)$ is introduced in Theorem 2.6. Combining (2.21) and [41, Theorem 2.6(2)], there exist constants $c_{1},\lambda_{1}>0$ independent of $\mu$ such that
@@
$\delta_{0}:=\left(\inf_{t>\frac{\log c_{1}}{\lambda_{1}}}\frac{\sqrt{\frac{1}{K_{1}}(\text{\rm{e}}^{K_{1}t}-1)}}{1-c_{1}\text{\rm{e}}^{-\lambda_{1}t}}\right)^{-2},$
By taking advantage of [45, Proposition 2.2] and ((H)), when $K_{I}<\min(K_{2},\delta_{0})$, (2.12) has a unique invariant probability measure $\mu_{\infty}\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$. For reader’s convenience, we give more details. For $\mu_{1},\mu_{2}\in\mathscr{P}_{2}(\mathbb{R}^{d})$, let $\tilde{X}_{t}^{i},i=1,2$ solve
@@
(3) Finally, assuming $\sigma(x,\gamma)=\sigma$ and (2.7), the log-Harnack inequality holds by [36], i.e.
@@
holds for some measurable function $\eta:(0,\infty)\rightarrow(0,\infty)$. Moreover, the Talagrand inequality also holds due to the log-Sobolev inequality (2.29), see for instance [27], which together with (2.3) and (2.30) implies (2.32), see [30] for more details. ∎
Remark 2.8.
Compared with the assumption $\sup_{\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})}|b(0,\mu)|<\infty$ in [45, Corollary 2.14], $|b(0,\mu)|$ can be of linear growth in $\mu$ in Theorem 2.7, see Example 2.9 below.
@@
$L^{\mu}=\langle b(\cdot,\mu),\nabla\rangle+\frac{1}{2}\mathrm{tr}((\sigma\sigma^{\ast})(\cdot,\mu)\nabla^{2}f),$
@@
then (2.2) is true, see for instance [2, 15]. Hence, the assertions in Theorem 2.6 and Theorem 2.7 also hold by replacing $\sigma(x,\gamma)=\sigma(\gamma)$ with (2.38). However, (2.38) is not easy to check.
@@ 3 Degenerate case
In this section, consider
@@
where $b:(0,\infty)\times\mathbb{R}^{2d}\times\mathscr{P}(\mathbb{R}^{2d})\rightarrow\mathbb{R}^{d}$, $\sigma:(0,\infty)\times\mathscr{P}(\mathbb{R}^{2d})\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{n}$ are measurable and bounded on bounded sets and $\{W_{t}\}_{t\geq 0}$ is an $n$-dimensional standard Brownian motion on some complete filtered probability space $(\Omega,\mathscr{F},(\mathscr{F}_{t})_{t\geq 0},\mathbb{P})$. As said above, when (3.3) is well-posed in $\mathscr{P}_{2}(\mathbb{R}^{2d})$, we use $P_{t}^{\ast}\mu_{0}$ to denote the distribution of the solution to (3.3) with initial distribution $\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$. For any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{2d}))$, we also consider the time-inhomogeneous decoupled SDEs:
@@
We make the following assumptions.
@@
    There exist constants $K_{M}>0,K_{I}>0,K>0$ such that
    $|b_{t}(z,\gamma)-b_{t}(\bar{z},\tilde{\gamma})|+\|\sigma_{t}(\gamma)-\sigma_{t}(\tilde{\gamma})\|_{HS}$
    $\leq K_{M}|z-\bar{z}|+K_{I}\mathbb{W}_{2}(\gamma,\tilde{\gamma}),\ \ z,\bar{z}\in\mathbb{R}^{2d},\gamma,\tilde{\gamma}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$
    and $|b_{t}(0,\delta_{0})|+\|\sigma_{t}(\delta_{0})\|_{HS}\leq K,\ \ t\geq 0.$
Under (B), (3.3) is well-posed in $\mathscr{P}_{2}(\mathbb{R}^{2d})$ and (3.6) is also well-posed. Let $P_{s,t}^{\mu}$ be the semigroup associated to (3.6). For a differentiable function $f$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$, $i=1,2$, let $\nabla^{(i)}f(x_{1},x_{2})$ denote the gradient of $f$ along $x_{i}$.
@@ 3.1 Log-Sobolev inequality for $P_{s,t}^{\mu}$ and $P_{t}^{\ast}\mu_{0}$
(3.7) $\sigma_{t}(\gamma)\sigma_{t}(\gamma)^{\ast}\leq\delta_{1},\ \ \gamma\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
Then the following assertions hold.
  1. (1)
    For any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{2d}))$, the log-Sobolev ineqiality holds: for $0\leq s\leq t,f>0,f\in C_{0}^{\infty}(\mathbb{R}^{2d})$
    (3.8) $P_{s,t}^{\mu}(f\log f)-P_{s,t}^{\mu}(f)\log P_{s,t}^{\mu}(f)\leq\frac{2\delta_{1}}{2K_{M}+1}(\text{\rm{e}}^{(2K_{M}+1)(t-s)}-1)P_{s,t}^{\mu}(|\nabla f^{\frac{1}{2}}|^{2}).$
  2. (2)
    If $\mu_{0}$ satisfies the log-Sobolev inequality: for some $L_{0}>0$,
    $\mu_{0}(f\log f)\leq\mu_{0}(f)\log\mu_{0}(f)+2L_{0}\mu_{0}(|\nabla f^{\frac{1}{2}}|^{2}),\ \ f>0,f\in C_{0}^{\infty}(\mathbb{R}^{2d})$
    then $P_{t}^{\ast}\mu_{0}$ also admits the log-Sobolev inequality
    $(P_{t}^{\ast}\mu_{0})(f\log f)-(P_{t}^{\ast}\mu_{0})(f)\log(P_{t}^{\ast}\mu_{0})(f)$
    $\leq 2\left\{\frac{\delta_{1}}{2K_{M}+1}(\text{\rm{e}}^{(2K_{M}+1)t}-1)+L_{0}\text{\rm{e}}^{(2K_{M}+1)t}\right\}(P_{t}^{\ast}\mu_{0})(|\nabla f^{\frac{1}{2}}|^{2}),$
    $\ \ f>0,f\in C_{0}^{\infty}(\mathbb{R}^{2d}),t\geq 0.$
Proof.
Let $\{X_{s,t}^{\mu,z}\}_{0\leq s\leq t}$ be the solution to (3.6) with $X_{s,s}^{\mu,z}=z\in\mathbb{R}^{2d}$. By (B), it is standard to derive that
$|X_{s,t}^{\mu,z}-X_{s,t}^{\mu,\bar{z}}|\leq\text{\rm{e}}^{(K_{M}+\frac{1}{2})(t-s)}|z-\bar{z}|,\ \ z,\bar{z}\in\mathbb{R}^{2d},0\leq s\leq t.$
This implies that
(3.9) $|\nabla P_{s,t}^{\mu}f|\leq\text{\rm{e}}^{(K_{M}+\frac{1}{2})(t-s)}P_{s,t}^{\mu}|\nabla f|,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d}),0\leq s\leq t.$
In the present case, we have $\Gamma^{t,\mu}_{1}(f)=|(\sigma_{t}^{\mu})^{\ast}\nabla^{(2)}f|^{2}$ and it is easy to see that
$\Gamma_{1}^{r,\mu}(P_{r,t}^{\mu}f)$ $=|(\sigma_{r}^{\mu})^{\ast}\nabla^{(2)}P_{r,t}^{\mu}f|^{2}\leq\delta_{1}|\nabla P_{r,t}^{\mu}f|^{2},\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d}),0\leq r\leq t.$
The remaining is completely the same with that of Theorem 2.1 with $K_{1}$ replaced by $12K_{M}+$ 1 and we omit it. ∎
@@
It is not difficult to see that (3.8) does not hold if $|\nabla f|^{2}$ is replaced by $|\nabla^{(2)}f|^{2}$. This is essentially different from the non-degenerate case. In fact, in the proof of Theorem 3.1, (3.9) does not hold if $\nabla$ is replaced by $\nabla^{(2)}$.
3.2 Hyperboundedness for time-homogenenous decoupled SDEs

### v4

@@ 2. $(2)$
Since $\sigma(x,\gamma)=\sigma(\gamma)$, by (2.6) with $s=0$, we have
@@
Since $\{\Phi(\mu):\mu\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})\}$ is tight due to (2.24), according to Lemma 2.10(2), in order to derive the desired log-Sobolev inequality (2.41), it remains to prove that for any $R>0$, there exist positive constants $C_{1}$ and $C_{2}$ depending on $R$, $M$ such that
@@
for $\rho_{\mu}:=\frac{\text{\rm{d}}\Phi(\mu)}{\text{\rm{d}}x}$. Since $b(x,\mu)$ is locally bounded in $(x,\mu)\in\mathbb{R}^{d}\times\mathscr{P}_{2}(\mathbb{R}^{d})$, for any $R>0$, there exists $\eta(R,M)>0$ such that
@@
Noting $(\sigma\sigma^{\ast})(\mu)\geq\delta_{2}I_{d\times d}$, by [5, Corollary 1.2.11], for any $R>0$, there exists a constant $C_{0}$ depending on $R$, $\eta(R,M)$ and $\delta_{2}$ such that
@@
2.3 Exponential ergodicity of McKean-Vlasov SDEs
We are now ready to prove exponential ergodicity of (2.20) in $L^{2}$-Wasserstein and relative entropy under the partially dissipative condition.
@@ 2.3 Exponential ergodicity of McKean-Vlasov SDEs
Assume (H) and that $\sigma(x,\gamma)=\sigma(\gamma)$.
  1. $(1)$
    If $K_{I}$ is small enough, then (2.20) has a unique IPM $\mu_{\infty}$ which satisfies the log-Sobolev inequality
    (2.44) $\mu_{\infty}(f\log f)-\mu_{\infty}(f)\log\mu_{\infty}(f)\leq C_{L}\mu_{\infty}(|\nabla f^{\frac{1}{2}}|^{2}),\ \ 0<f\in C_{0}^{\infty}(\mathbb{R}^{d})$
    for some constant $C_{L}>0$.
  2. $(2)$
    Moreover, there exist constants $c>0,\lambda>0$ such that
@@ 3. $(3)$
    (2.47) $\mathrm{Ent}(P_{t}^{\ast}\mu_{0}|\mu_{\infty})\leq\bar{c}\text{\rm{e}}^{-2\lambda t}\mathrm{Ent}(\mu_{0}|\mu_{\infty}),\ \ t\geq 1,\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{d})$
@@
Remark 2.14.
As explained in the introduction, to derive the uniform in time log-Sobolev inequality, [35, Proposition 3.9] requires .
$\left\langle\nabla U(x)+\nabla_{x}W(x,z)-\nabla U(y)-\nabla_{y}W(y,z),x-y\right\rangle\leq-\rho|x-y|^{2},\ \ |x|\vee|y|\geq R,$
which implies
$-\left\langle\nabla U(x)+\int_{\mathbb{R}^{d}}\nabla_{x}W(x,z)\mu(\text{\rm{d}}z)-\nabla U(y)-\int_{\mathbb{R}^{d}}\nabla_{y}W(y,z)\mu(\text{\rm{d}}z),x-y\right\rangle$
$\leq-\rho|x-y|^{2},\ \ |x-y|\geq 2R.$
So, the condition on the drift in [35, Proposition 3.9] is slightly stronger than (H). Besides, we do not require the high diffusivity as in [35, Proposition 3.9]. Moreover, the assumption on $\sigma$ in Theorem 2.13(3) is to derive the log-Harnack inequality. To our best knowledge, the log-Harnack inequality for McKean-Vlasov SDEs with $\sigma$ being dependent on the measure variable and $b$ satisfying ($(3)$) is still open.
Remark 2.15.
Compared with the assumption $\sup_{\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})}|b(0,\mu)|<\infty$ in [56, Corollary 2.14], $|b(0,\mu)|$ can be of linear growth in $\mu$ in Theorem 2.13, see Example 2.16 below. The condition of $K_{I}$ being small enough is essential, otherwise the uniqueness of invariant probability measures may fail, see for instance [14].
@@
(1) Recall $\Phi(\mu)$ is given in Theorem 2.6. Combining (2.41) and [52, Theorem 2.6(2)], there exist constants $c_{1},\lambda_{1}>0$ independent of $\mu$ such that
@@
$\delta_{0}:=\left(\inf_{t>\frac{\log c_{1}}{\lambda_{1}}}\frac{\sqrt{\frac{1}{K_{1}}(\text{\rm{e}}^{K_{1}t}-1)}}{1-c_{1}\text{\rm{e}}^{-\lambda_{1}t}}\right)^{-2}.$
By taking advantage of [56, Proposition 2.2] and ((H)), when $K_{I}<\min(K_{2},\delta_{0})$, (2.20) has a unique invariant probability measure $\mu_{\infty}\in\hat{\mathscr{P}}_{2,M}(\mathbb{R}^{d})$. For reader’s convenience, we give more details. For $\mu_{1},\mu_{2}\in\mathscr{P}_{2}(\mathbb{R}^{d})$, let $\tilde{X}_{t}^{i},i=1,2$ solve
@@
(3) Finally, assuming $\sigma(x,\gamma)=\sigma$ and ($(3)$), the log-Harnack inequality holds by [47], i.e.
@@
holds for some measurable function $\eta:(0,\infty)\rightarrow(0,\infty)$. Moreover, the Talagrand inequality also holds due to the log-Sobolev inequality (2.44), see for instance [37], which together with (2.3) and (2.45) implies (2.47), see for instance [41] for more details. ∎
@@
$L^{\mu}=\langle b(\cdot,\mu),\nabla\rangle+\frac{1}{2}\mathrm{tr}((\sigma\sigma^{\ast})(\cdot,\mu)\nabla^{2}),$
@@
then (2.2.3) is true, see for instance [2, 23]. Hence, the assertions in Theorem 2.11 and Theorem 2.13 also hold by replacing $\sigma(x,\gamma)=\sigma(\gamma)$ with (2.53). However, (2.53) is not easy to check.
@@ 3 Degenerate case
In this section, we consider
@@
where $b:[0,\infty)\times\mathbb{R}^{2d}\times\mathscr{P}(\mathbb{R}^{2d})\rightarrow\mathbb{R}^{d}$, $\sigma:[0,\infty)\times\mathscr{P}(\mathbb{R}^{2d})\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{n}$ are measurable and bounded on bounded sets and $\{W_{t}\}_{t\geq 0}$ is an $n$-dimensional standard Brownian motion on some complete filtered probability space $(\Omega,\mathscr{F},(\mathscr{F}_{t})_{t\geq 0},\mathbb{P})$. As said above, when (3.3) is well-posed in $\mathscr{P}_{2}(\mathbb{R}^{2d})$, we use $P_{t}^{\ast}\mu_{0}$ to denote the distribution of the solution to (3.3) with initial distribution $\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$. For any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{2d}))$, we also consider the time-inhomogeneous decoupled SDEs:
@@
We make the following assumptions
@@
    $b$ is bounded on bounded sets of $0,\infty)\times\mathbb{R}^{2d}\times\mathscr{P}_{2}(\mathbb{R}^{2d})$. For any $t\geq 0$ and $\gamma\in\mathscr{P}_{2}(\mathbb{R}^{2d})$, $b_{t}(\cdot,\gamma)$ is continuous in $\mathbb{R}^{2d}$. There exist constants $K_{1},\tilde{K}_{I}>0$ such that for all $t\geq 0$, $x,\tilde{x},y,\tilde{y}\in\mathbb{R}^{d}$ and $\gamma,\tilde{\gamma}\in\mathscr{P}_{2}(\mathbb{R}^{2d})$,
    $2\langle b_{t}(x,y,\gamma)-b_{t}(\tilde{x},\tilde{y},\tilde{\gamma}),y-\tilde{y}\rangle+2\langle y-\tilde{y},x-\tilde{x}\rangle+\|\sigma_{t}(\gamma)-\sigma_{t}(\tilde{\gamma})\|_{HS}^{2}$
    $\leq K_{1}(|x-\tilde{x}|^{2}+|y-\tilde{y}|^{2})+\tilde{K}_{I}\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},$
    and
    $\|\sigma_{t}(\gamma)\|_{HS}^{2}\leq K_{1}(1+\gamma(|\cdot|^{2})).$
Let
$B_{t}(x,y,\gamma)=\left(\begin{array}[]{c}y\\ b_{t}(x,y,\gamma)\\ \end{array}\right),\ \ \Sigma_{t}(\gamma)=\left(\begin{array}[]{c}0_{d\times n}\\ \sigma_{t}(\gamma)\\ \end{array}\right),\ \ x,y\in\mathbb{R}^{d},\gamma\in\mathscr{P}(\mathbb{R}^{2d}).$
Let $Z_{t}=\left(\begin{array}[]{c}X_{t}\\ Y_{t}\\ \end{array}\right)$ and $Z_{t}^{\mu}=\left(\begin{array}[]{c}X_{t}^{\mu}\\ Y_{t}^{\mu}\\ \end{array}\right)$. Then (3.3) and (3.6) can be rewritten as
$\text{\rm{d}}Z_{t}=B_{t}(Z_{t},\mathscr{L}_{Z_{t}})+\Sigma_{t}(\mathscr{L}_{Z_{t}})\text{\rm{d}}W_{t}$
and
$\text{\rm{d}}Z_{t}^{\mu}=B_{t}(Z_{t}^{\mu},\mu_{t})+\Sigma_{t}(\mu_{t})\text{\rm{d}}W_{t}.$
So, under (B), Theorem 4.1 implies that (3.3) is well-posed in $\mathscr{P}_{2}(\mathbb{R}^{2d})$ and (3.6) is also well-posed. Let $P_{s,t}^{\mu}$ be the semigroup associated to (3.6).
For a differentiable function $f$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$, $i=1,2$, let $\nabla^{(i)}f(x_{1},x_{2})$ denote the gradient of $f$ along $x_{i}$.
@@ 3.1 Log-Sobolev inequality for $P_{s,t}^{\mu}$ and $P_{t}^{\ast}\mu_{0}$
(3.7) $\sigma_{t}(\gamma)\sigma_{t}(\gamma)^{\ast}\leq\delta_{1}I_{d\times d},\ \ \gamma\in\mathscr{P}_{2}(\mathbb{R}^{2d}),t\geq 0.$
Then the assertions in Theorem 2.1 hold for (3.3) and (3.6) on $\mathbb{R}^{2d}$ replacing (1.1) and (2.1) on $\mathbb{R}^{d}$.
@@
It is not difficult to see that the log-Sobolev inequality (2.6) does not hold if $\nabla$ is replaced by $\nabla^{(2)}$. This is essentially different from the non-degenerate case. In fact, the inequality
$|\nabla P_{s,t}^{\mu}f|\leq\text{\rm{e}}^{\frac{K_{1}}{2}(t-s)}P_{s,t}^{\mu}|\nabla f|,\ \ f\in C_{b}^{\infty}(\mathbb{R}^{2d}),0\leq s\leq t$
does not hold if $\nabla$ is replaced by $\nabla^{(2)}$.
3.2 Time-homogeneous decoupled SDEs